OpenStudy (anonymous):

MEDAL... A certain paint is to be mixed with 3 parts yellow to 2 parts blue. Twelve gallons of mixed paint are needed. How many gallons of blue must be used? 4 4.8 7.2 10

OpenStudy (anonymous):

@ash2326 @electrokid help?

OpenStudy (anonymous):

4.8

OpenStudy (anonymous):

u need explanation @iribeiro1602

OpenStudy (gorv):

ratio of yellow and blue is 3:2 that is in 5 gallon paint 3/5 part is yellow and 2/5 part is blue so in 12 gallon blue=(2/5)*12 =4.8

OpenStudy (anonymous):

\[ \frac{\rm yellow}{\rm blue}={3\over2}\\ {\rm yellow}+{\rm blue}=12\\ {\rm yellow}+{3\over2}{\rm yellow}=12\\ \frac{2{\rm yellow}+3{\rm yellow}}{2}=12\\ 5\times{\rm yellow}=12\times 2\\ {\rm yellow}={12\times 2\over5}=? \]

OpenStudy (anonymous):

now, that we have the amount of "yellow", the amount of "blue"=12-"amount of yellow"=?